My son is back in school, and my algebra from 20+ years ago is not coming back to me...
Please help me with this:
Julie is 4 times older than Aaron. In 10 years Julie will be twice as old as Aaron. How old are Julie and Aaron?

Answer by stanbon(75887)

Julie is 4 times older than Aaron. In 10 years Julie will be twice as old as Aaron. How old are Julie and Aaron?
------------------------------------------
Now DATA:
Let Aaron be x yrs old.
Then Julie is 4x yrs old
-------------------------
10 years from now DATA:
Aaron will be x+10 yrs old.
Julie will be 4x+10 yrs old.
------------------------------
EQUATION:
Julie = twice Aaron's age.
4x+10 = 2(x+10)
4x + 10 = 2x + 20
2x = 10
x = 5 (Aaron's age NOW)
4x = 20 (Julie's age NOW)
=============================
Cheers,
Stan H.